OpenStudy (anonymous):

solving for the x-intercept of 1=2^x

myininaya (myininaya):

is how the question looks exactly?

OpenStudy (anonymous):

g(x)=2^x-1

myininaya (myininaya):

ok to find x-intercept set y=0 and solve for x (y=g(x))

myininaya (myininaya):

0=2^x-1 1=2^x ln1=ln2^x 0=xln2 x=0

OpenStudy (anonymous):

x-intercept means y=0, or find g(0). g(0)=2^0-1 g(0)=1-1 g(0)=0 x-intercept is 0...

myininaya (myininaya):

x intercept is (0,0) this is also a y-intecept

OpenStudy (anonymous):

thank you